# Learning parameter dependence for Fourier-based option pricing with tensor trains

### v6

@@ VI Results
The results are summarized in TABLE 2. In particular, the computational complexity versus $d$ is plotted in FIG. 2. The maximum error in TT-based option prices among runs for 100 random parameter sets is represented by $e_{\mathrm{TT}}$. For the parameter set that yields this maximum error, we compute the half-width of the 95% confidence interval for the MC-based option price and denote it by $e_{\mathrm{MC},10^{5}}$ and $e_{\mathrm{MC},10^{6}}$ for the cases with $10^{5}$ and $10^{6}$ paths, respectively. The computational complexity and time of TT-based option pricing are represented by $c_{\mathrm{TT}}$ and $t_{\mathrm{TT}}$, respectively, and those of MC-based pricing with $10^{5}$ paths are denoted by $c_{\mathrm{MC},10^{5}}$ and $t_{\mathrm{MC},10^{5}}$, with those for $10^{6}$ paths being 10 times larger. To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$ for $d\leq 10$, and $\epsilon_{\mathrm{TCI}}=10^{-10}$ for $d=11$. Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively. The maximum bond dimensions of the TTs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2. In the SVD on $\tilde{V}_{k_{1},\ldots,k_{d}}$ obtained by contracting the TTs for $\phi$ and $\hat{v}$, we set the tolerance $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$, and the maximum bond dimension of the resulting TT is denoted by $\chi_{V}$.
@@ VI.0.1 The case of varying $\vec{\sigma}$
TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence. TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method. The bond dimensions of the TT results for $d=5$ and $d=10$ are depicted in FIG. 3(a). By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at around 10 when $\eta$ is constant. When $\eta$ is automatically determined for each asset, the bond dimensions are around 60 or 70. The details of this compression by SVD are described in Appendix. The maximum bond dimension $\chi_{\tilde{V}}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is maintained at 2, regardless of when $\eta$ is constant or $\eta$ is automatically determined for each asset. Therefore, the computational complexity is much lower than that of MC-based option pricing.
We compare the TT-based results with the MC simulations with $10^{5}$ and $10^{6}$ paths. With $10^{5}$ paths, the TT-based option price is within the MC error range in all the tested cases. With $10^{6}$ paths, the TT-based method overall gives the errors comparable with the MC-based method, though it performes worse for $d=8$ and $d=9$.
Furthermore, reducing the number of subdivisions below 100 resulted in only minor changes in accuracy and computational complexity.
@@
$d$ $\eta$ $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ $e_{\mathrm{TT}}$ $e_{\text{MC},10^{5}}$ $e_{\text{MC},10^{6}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC},10^{5}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC},10^{5}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
$5$ 0.9 $1.0\times 10^{-5}$ 0.00560 0.0213 0.00670 $14$ $5.0\times 10^{5}$ $7.77\times 10^{-7}$ $2.84\times 10^{-2}$ 12 9 2
$6$ 0.75 $1.0\times 10^{-5}$ 0.00377 0.0143 0.00456 $18$ $6.0\times 10^{5}$ $9.37\times 10^{-7}$ $3.38\times 10^{-2}$ 13 9 2
$7$ 0.69 $1.0\times 10^{-5}$ 0.00380 0.0124 0.00382 $22$ $7.0\times 10^{5}$ $9.12\times 10^{-7}$ $3.71\times 10^{-2}$ 13 8 2
$8$ 0.66 $1.5\times 10^{-5}$ 0.00547 0.0128 0.00346 $26$ $8.0\times 10^{5}$ $1.75\times 10^{-6}$ $4.09\times 10^{-2}$ 14 8 2
$9$ 0.61 $1.5\times 10^{-5}$ 0.00694 0.0100 0.00302 $30$ $9.0\times 10^{5}$ $2.47\times 10^{-6}$ $4.51\times 10^{-2}$ 13 7 2
$10$ auto $1.5\times 10^{-6}$ 0.00230 0.00916 0.00282 $34$ $1.0\times 10^{6}$ $2.00\times 10^{-6}$ $5.11\times 10^{-2}$ 59 58 2
$11$ auto $1.5\times 10^{-7}$ 0.00117 0.00847 0.00269 $38$ $1.1\times 10^{7}$ $2.16\times 10^{-6}$ $5.04\times 10^{-2}$ 73 76 2
@@
$d$ $\eta$ $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ $e_{\mathrm{TT}}$ $e_{\text{MC},10^{5}}$ $e_{\text{MC},10^{6}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC},10^{5}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC},10^{5}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
$5$ 0.9 $1.0\times 10^{-5}$ 0.0195 0.0420 0.0133 $14$ $5.0\times 10^{5}$ $7.67\times 10^{-7}$ $3.99\times 10^{-2}$ 14 9 2
$6$ 0.75 $1.0\times 10^{-5}$ 0.0165 0.0385 0.0121 $18$ $6.0\times 10^{5}$ $9.34\times 10^{-7}$ $5.15\times 10^{-2}$ 16 9 2
$7$ 0.69 $3.0\times 10^{-5}$ 0.0103 0.0357 0.0112 $22$ $7.0\times 10^{5}$ $9.21\times 10^{-7}$ $5.37\times 10^{-2}$ $21$ $10$ $2$
$8$ 0.57 $3.5\times 10^{-5}$ 0.0125 0.0303 0.00952 $26$ $8.0\times 10^{5}$ $1.63\times 10^{-6}$ $6.18\times 10^{-2}$ $17$ $7$ $2$
$9$ 0.53 $1.0\times 10^{-5}$ 0.0155 0.0204 0.00650 $30$ $9.0\times 10^{5}$ $1.19\times 10^{-6}$ $7.95\times 10^{-2}$ $19$ $7$ $2$
$10$ 0.5 $3.0\times 10^{-5}$ 0.0172 0.0204 0.00654 $34$ $1.0\times 10^{6}$ $1.47\times 10^{-6}$ $8.77\times 10^{-2}$ $16$ $6$ $2$
$11$ auto $1.0\times 10^{-7}$ 0.0139 0.0181 0.00579 $38$ $1.1\times 10^{6}$ $1.93\times 10^{-6}$ $8.43\times 10^{-2}$ $48$ $38$ $2$
Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence. Here, we set the ranges $\sigma_{m}\in[0.15,0.25]$ and $S_{m,0}\in[90,120]$, and place 100 equally spaced grid points within these ranges. The $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ in the table refers to the values adjusted so that $c_{\mathrm{TT}}$ is reduced with $e_{\mathrm{TT}}$ kept smaller than the MC error $e_{\text{MC},10^{5}}$. The term auto in (a) and (b) indicate that $\eta_{m}$ was automatically determined for each asset based on Eqs. (19) and (20), respectively. The computational complexity and time increases when $\eta$ is set by Eq. (19) or (20). $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$ versus the number of assets $d$. We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs and the number of path in MC are $10^{5}$ and $10^{6}$.
Figure 3: The bond dimensions $\chi_{l}$ of each bond $l$ for $\phi$ and $\hat{v}$, obtained through TCI and SVD, incorporating dependencies on (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$. The odd bond $l=2i-1$ connects the core tensors for $j_{m}$ and $k_{m}$, the indexes for $z_{m}$ and $p_{m}$, respectively, and the even bond $l=2i$ connects those for $k_{m}$ and $j_{m+1}$. In the cases corresponding to the left panels and the bottom right panel, $\hat{v}$ does not depend on parameters, and thus the bond $l$ takes only even values $2i$ (connecting the core tensors for $z_{m}$ and $z_{m+1}$). It is noteworthy that the graph of bond dimensions exhibits a characteristic jagged shape. The bond dimensions between sites for $j_{m}$ and $k_{m}$, which are related to the same asset, are large, and the ones between sites for $k_{m}$ and $j_{m+1}$, which are related to different assets, are small.
@@ VI.0.2 The case of varying $\vec{S_{0}}$
TABLE 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence. TT-based option pricing demonstrates superiority over the Monte Carlo method in terms of computational complexity and time. For $d$=5 and 10, after compression using SVD and contractions, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ were reduced to around 10 (refer to FIG. 3(b)). For $d=11$, when $\eta$ is determined automatically, this bond dimension increases to around 40. The maximum bond dimension, $\chi_{\tilde{V}}$, of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is 2 Therefore, we saw superiority over the Monte Carlo method in terms of computation complexity and time.
Compared with the MC-based method with $10^{5}$ paths, the TT-based mathod again gives the smaller error in every value of $d$. With $10^{6}$ paths, the TT error is worse than the MC error for many values of $d$, but still comparable.
@@ VI.0.3 Randomness in learning the TTs
Here, we mention the randomness in learning the TTs and the error induced by it. Since TCI starts with the randomly selected grid points, the accuracy of the TT-based method depends on these initial points. To assess such a fluctuation of the accuracy, for $d=10$, we evaluated the mean and standard deviation of the relative accuracy of the TT-based method in 20 runs, with initial points randomly selected in each. As a result, the mean was 0.0324 and the standard deviation was 0.0039, which indicates that the accuracy fluctuation of TT-based option pricing is very small. We should note that TCI is a heuristic method, and depending on the choice of the initial points, the learning might not work well. That is, the error defined by Eq. (3) might not go below the threshold. Here, with the tolerance for TCI/SVD fixed, if the maximum relative error exceeds 4%, we redo TCI and SVD. The aforementioned mean and standard deviation are computed from successful cases.
@@ VI.0.4 Total computational time for obtaining the TTs
We mention the total computational time for obtaining the TTs for $\phi$ and $\hat{v}$ and $\tilde{V}_{k_{1},\ldots,k_{d}}$ through TCI and SVD. We focus on TCI because it dominates over SVD. It took about 3.4 hours for $d=11$ in the case that $\vec{\sigma}$ dependence is involved. This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec. VII. We also note that, when we do not incorporate the parameter dependence into the tensor trains for $\phi$ and $\hat{v}$ as in Ref. [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence. In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$ and all other parameters are the same as those in the default settings listed in Table 1. This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
@@ VII Discussion
We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from functions of Eqs. (10) and (12), enabling fast option pricing in response to varying parameters. In this study, we considered scenarios with varying volatility and present stock prices as benchmarks for our proposed method. Up to $d=11$, we demonstrated superiority in both computational complexity and time. Note that in the MC-based method, the implementation may have room for improvement to reduce computational time.
We numerically verified that all TT-based results fell within the statistical error range of the MC-based method with $10^{5}$ paths. Compared to the MC-based method with $10^{6}$ paths, the TT-based method gave the worse errors in some cases, but was still comparable. To enhance the accuracy of the TT-based method, we may reduce the tolerance of the TCI further or change the initial randomly selected grid points of TCI.
@@
Now, let us consider how the proposed method provides benefits in practical business in financial firms. An expected way to utilize this method is as follows. At night, when the financial market is closed, we learn the TTs and perform contractions, and then, during the day, when the market is open, we use the TT to quickly price the option for the fluctuating parameters. If we pursue the computational speed in the daytime and allow the overnight precomputation to some extent, the above operation can be beneficial. In light of this, it is reasonable that we compare the computational complexity in the TT-based method after the TTs are obtained with that in the MC-based method, neglecting the learning process and contraction.
@@
Finally, we discuss future research directions. If we could keep $\eta$ constant when $d$ is increased, we have the advantage of reducing the computational complexity involved in option pricing. It may be possible to keep $\eta$ constant by adjusting the value of parameter $N$, e.g. by increasing $N$ from the current value of 50 to, say, 100. It could be a promising approach to directly learning the TTOs incorporating parameter dependence using TCI, potentially improving the accuracy of option pricing. At the same time, it is important to study methods to improve the numerical stability and accuracy of TCI itself. For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TN format is desired, although this study tried to incorporate either of them as the first step. Besides, to enhance the practical benefit, we should extend our methodology so that it is applicable to more advanced settings in option pricing. With respect to the pricing model, while this paper has considered the BS model, more advanced models such as the local volatility model [38] and stochastic volatility model [39, 40] are also used in practice. Applying our method to such models, where the dependencies on the parameters in the models are incorporated into TTs, is an interesting direction. Expanding the scope to the broad types of products is also desired. While this paper focused on European-type options, more complicated ones such as American and Bermudan options are also traded widely, and pricing of such options is more time-consuming. Considering whether our method can be extended for such options will be an important and interesting challenge.
VIII Acknowledgements

### v7

@@ VI Results
The results are summarized in TABLE 2. In particular, the computational complexity versus $d$ is plotted in FIG. 2. The mean absolute error in TT-based option prices among runs for 100 random parameter sets is represented by $e_{\mathrm{TT}}$. For the same parameter sets, we compute mean values of the absolute values of the error in the MC-based option price with $10^{6}$ paths and denote it by $e_{\mathrm{MC}}$. The computational complexity and time of TT-based option pricing are represented by $c_{\mathrm{TT}}$ and $t_{\mathrm{TT}}$, respectively, and those of MC-based pricing with $10^{6}$ paths are denoted by $c_{\mathrm{MC}}$ and $t_{\mathrm{MC}}$. To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$. Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}=1.0\times 10^{-6}$ for $\phi$ and $\hat{v}$. The maximum bond dimensions of the TTs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2. In the SVD on $\tilde{V}_{k_{1},\ldots,k_{d}}$ obtained by contracting the TTs for $\phi$ and $\hat{v}$, we set the tolerance $\epsilon^{\tilde{V}}_{\mathrm{SVD}}=1.0\times 10^{-6}$, and the maximum bond dimension of the resulting TT is denoted by $\chi_{V}$.
@@ VI.0.1 The case of varying $\vec{\sigma}$
TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence. TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method. The bond dimensions of the TT results for $d=5$ and $d=10$ are depicted in FIG. 3(a). By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at around 10 when $\eta$ is constant. The details of this compression by SVD are described in Appendix. A. The maximum bond dimension $\chi_{\tilde{V}}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is maintained at 2, or 1 especially when $d=10,11$. Therefore, the computational complexity is much lower than that of MC-based option pricing. Besides, the error in the TT-based result is smaller than that in the MC-based result in all the tested cases.
@@
$d$ $e_{\mathrm{TT}}$ $e_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
$5$ 0.00178 0.00606 $16$ $5.0\times 10^{6}$ $4.91\times 10^{-7}$ $2.84\times 10^{-1}$ 16 11 2
$6$ 0.00154 0.00503 $20$ $6.0\times 10^{6}$ $6.37\times 10^{-7}$ $3.38\times 10^{-1}$ 16 11 2
$7$ 0.00134 0.00428 $24$ $7.0\times 10^{6}$ $9.12\times 10^{-7}$ $3.71\times 10^{-1}$ 17 11 2
$8$ 0.00136 0.00372 $28$ $8.0\times 10^{6}$ $1.75\times 10^{-6}$ $4.09\times 10^{-1}$ 18 11 2
$9$ 0.000867 0.00329 $32$ $9.0\times 10^{6}$ $2.47\times 10^{-6}$ $4.51\times 10^{-1}$ 20 11 2
$10$ 0.00229 0.00294 $10$ $1.0\times 10^{7}$ $1.18\times 10^{-6}$ $5.11\times 10^{-1}$ 20 11 1
$11$ 0.000554 0.00265 $11$ $1.1\times 10^{7}$ $1.16\times 10^{-6}$ $5.04\times 10^{-1}$ 20 11 1
@@
$d$ $e_{\mathrm{TT}}$ $e_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
$5$ 0.00151 0.00773 $16$ $5.0\times 10^{6}$ $7.67\times 10^{-7}$ $3.99\times 10^{-1}$ 18 11 2
$6$ 0.00122 0.00640 $20$ $6.0\times 10^{6}$ $6.34\times 10^{-7}$ $5.15\times 10^{-1}$ 19 11 2
$7$ 0.00112 0.00547 $24$ $7.0\times 10^{6}$ $9.21\times 10^{-7}$ $5.37\times 10^{-1}$ $21$ $10$ $2$
$8$ 0.000973 0.00477 $28$ $8.0\times 10^{6}$ $1.63\times 10^{-6}$ $6.18\times 10^{-1}$ $23$ $11$ $2$
$9$ 0.000686 0.00424 $32$ $9.0\times 10^{6}$ $1.19\times 10^{-6}$ $7.95\times 10^{-1}$ $24$ $11$ $2$
$10$ 0.000662 0.00377 $36$ $1.0\times 10^{7}$ $1.47\times 10^{-6}$ $8.77\times 10^{-1}$ $24$ $11$ $2$
$11$ 0.00114 0.00339 $40$ $1.1\times 10^{7}$ $1.93\times 10^{-6}$ $8.43\times 10^{-1}$ $25$ $13$ $2$
Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence. Here, we set the ranges $\sigma_{m}\in[0.15,0.25)$ and $S_{m,0}\in[90,120)$, and place 100 equally spaced grid points within these ranges. $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$ versus the number of assets $d$. We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs and the number of path in MC are $10^{6}$.
Figure 3: The bond dimensions $\chi_{l}$ of each bond $l$ for $\phi$ and $\hat{v}$, obtained through TCI and SVD, incorporating dependencies on (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$. The odd bond $l=2i-1$ connects the core tensors for $j_{m}$ and $k_{m}$, the indexes for $z_{m}$ and $p_{m}$, respectively, and the even bond $l=2i$ connects those for $k_{m}$ and $j_{m+1}$. Note that $\hat{v}$ does not depend on parameters, and thus the bond $l$ takes only even values $2i$ (connecting the core tensors for $z_{m}$ and $z_{m+1}$). It is noteworthy that the graph of bond dimensions exhibits a characteristic jagged shape. The bond dimensions between sites for $j_{m}$ and $k_{m}$, which are related to the same asset, are large, and the ones between sites for $k_{m}$ and $j_{m+1}$, which are related to different assets, are small.
@@ VI.0.2 The case of varying $\vec{S_{0}}$
TABLE 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence. TT-based option pricing demonstrates superiority over the Monte Carlo method in terms of computational complexity and time. For $d$=5 and 10, after compression using SVD and contractions, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ were reduced to around 10 (refer to FIG. 3(b)). The maximum bond dimension, $\chi_{\tilde{V}}$, of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is 2. Therefore, we saw superiority over the Monte Carlo method in terms of computation complexity and time.
Compared with the MC-based method with $10^{6}$ paths, the TT-based mathod again gives the smaller error in every value of $d$.
@@ VI.0.3 Randomness in learning the TTs
Here, we mention the randomness in learning the TTs and the error induced by it. We should note that TCI is a heuristic method, and depending on the choice of the initial points, the learning might not work well. That is, the error defined by Eq. (3) might not go below the threshold. To assess such a fluctuation of the accuracy, for $d=10$ in the case of varying $\vec{\sigma}$, we evaluated the mean and standard deviation of the accuracy of the TT-based method in 20 runs, with initial points randomly selected in each. As a result, the mean was 0.00230 and the standard deviation was $4.74\times 10^{-9}$, which indicates that the accuracy fluctuation of TT-based option pricing is very small.
@@ VI.0.4 Total computational time for obtaining the TTs
We mention the total computational time for obtaining the TTs for $\phi$ and $\hat{v}$ and $\tilde{V}_{k_{1},\ldots,k_{d}}$ through TCI and SVD. We focus on TCI because it dominates over SVD. It took about $12.5$ minutes for $d=11$ in the case that $\vec{S}_{0}$ dependence is involved. This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec. VII. We also note that, when we do not incorporate the parameter dependence into the tensor train for $\phi$ as in Ref. [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence. In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$, $N=50$ and all other parameters are the same as those in the default settings listed in Table 1. This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
@@ VII Discussion
We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from the function of Eq. (10), enabling fast option pricing in response to varying parameters. In this study, we considered scenarios with varying volatility and present stock prices as benchmarks for our proposed method. Up to $d=11$, we demonstrated superiority in both computational complexity and time. Note that in the MC-based method, the implementation may have room for improvement to reduce computational complexity and time. We have also seen that in all the tested cases, the error in the TT-based result is smaller than that of the MC-based result with $10^{6}$ paths.
@@
Now, let us consider how the proposed method provides benefits in practical business in financial firms. An expected way to utilize this method is as follows. At night, when the financial market is closed, we learn the TTs and perform contractions, and then, during the day, when the market is open, we use the TT of parameter-dependent option prices to quickly price the option for the fluctuating parameters. If we pursue the computational speed in the daytime and allow the overnight precomputation to some extent, the above operation can be beneficial. In light of this, it is reasonable that we compare the computational complexity in the TT-based method after the TTs are obtained with that in the MC-based method, neglecting the learning process and contraction.
@@
Finally, we discuss future research directions. For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TN format is desired, although this study tried to incorporate either of them as the first step. Besides, to enhance the practical benefit, we should extend our methodology so that it is applicable to more advanced settings in option pricing. With respect to the pricing model, while this paper has considered the BS model, more advanced models such as the local volatility model [39], stochastic volatility model [40, 41] and Lévy models [42, 43] are also used in practice. Applying our method to such models, where the dependencies on the parameters in the models are incorporated into TTs, is an interesting direction. Expanding the scope to the broad types of products is also desired. While this paper focused on European-type options, more complicated ones such as American and Bermudan options are also traded widely, and pricing of such options is more time-consuming. Considering whether our method can be extended for such options will be an important and interesting challenge.
Acknowledgements